# Phase error rate estimation in QKD with imperfect detectors

### v2

@@ I Introduction
Security proofs of QKD based on the entropic uncertainty relations (EUR) [1, 2, 3], and the phase error correction approach [4, 5], yield some of the highest key rates against coherent attacks in the finite-size regime. While source imperfections [6, 7, 8, 9] have been extensively studied within the phase error correction framework, detector imperfections have not yet been addressed in any meaningful sense in either security proof framework. In particular, these proof techniques require the probability of detection in Bob’s detection setup to be independent of basis choice. This assumption is referred to as “basis-independent loss” in the literature, while the violation of this assumption is referred to as “basis-efficiency mismatch” or “detection-efficiency mismatch”. Satisfying this assumption requires the efficiency and dark count rates of Bob’s detectors to be exactly identical. Therefore, justifying this assumption in practice requires exact characterization of Bob’s identical detectors. Either of these tasks are impossible in practice. Therefore, these proof techniques are not applicable to practical prepare-and-measure (and entangled-based) QKD scenarios involving realistic detectors. This has led to many works advocating for the use of MDI-QKD [10] to address detector imperfections. To date, there is no security proof method that addresses this problem without requiring significant protocol modifications (such as MDI-QKD). While some theoretical analyses of basis-efficiency mismatch in the IID asymptotic setting exist for standard QKD [11, 12, 13, 14, 15], there is no work that handles basis-efficiency mismatch for coherent attacks in the finite-size regime. Meanwhile, there have been several experimental demonstrations [16, 17, 18, 19, 20], exploiting basis-efficiency mismatch for attacks on QKD systems.
In this work, we prove the security of the decoy-state BB84 protocol with an active detection setup without assuming basis-independent loss. We do so by showing that the phase error rate can be suitably bounded even without the assumption. We explicitly define metrics $\delta_{1}$, $\delta_{2}$ that quantify the deviation from the ideal case, and bound the phase error rate in terms of these deviations. We compute these metrics for the case of detectors with basis-efficiency mismatch, directly from the experimental characterization of the detectors. Our results extend the security of QKD to the following practical scenarios:
@@
    Bob’s detectors are not identical, but the values of efficiency ($\eta_{b_{i}}$ for basis $b$ and outcome $i$) and dark count rates ($d_{b_{i}}$) are known. Note that while this is a useful toy model, such scenarios are impractical since they require $\eta_{b_{i}},d_{b_{i}}$ to be known exactly.
@@
    Bob’s detectors are not identical, and the values of efficiency and dark count rates are only known to be in some range. Moreover, these values depend on the spatio-temporal modes (labelled by $d$) of the incoming photons, and can therefore be controlled by Eve [17, 13, 19, 20]. This is expressed mathematically as $\eta_{b_{i}}(d)\in[\eta_{\text{det}}(d)(1-\Delta_{\eta}),\eta_{\text{det}}(d)(1+\Delta_{\eta})],d_{b_{i}}(d)\in[d_{\text{det}}(1-\Delta_{\text{dc}}),d_{\text{det}}(1+\Delta_{\text{dc}})]$. Note that in this model, the range of allowed values of the loss can depend on the spatio-temporal mode, whereas the dark count rates for all the modes lie in the same range. This is a very realistic scenario.
@@
Our metrics $\delta_{1},\delta_{2}$ involve an optimization over all possible values of $\eta_{b_{i}},d_{b_{i}}$ in their respective ranges. Moreover, we find that our methods yield the same values for Case 2 and Case 3. Thus, the practically important problem of detector side-channels is a by-product of our proof technique.
@@
We use entropic uncertainty relations [1] for our security proofs in this work. Since both the entropic uncertainty relations approach and the phase error correction [4, 5] approach involve bounding the phase error rate, we expect our techniques to also work when using the phase error correction approach for the security proof. In fact, one may also use the equivalence by Tsurumuru [21, 22] to relate the two security proof approaches. For example, [22, Section III] constructs an explicit phase error correction circuit whose failure probability can be bound by the smooth min entropy via [22, Theorem 1 and Corollary 2]. Therefore one may indirectly prove security in the phase error correction framework by applying the above equivalence on the results of this work.
@@
Moreover, we prove security in the variable-length framework [23] which allows Alice and Bob to adaptively determine the number of bits to be used for error-correction, and the length of the output key, based on the observations during runtime. Such protocols are critical for practical implementations, where the honest behaviour of the channel connecting Alice and Bob is difficult to determine in advance. A security proof for variable-length QKD has been recently obtained for generic protocols in Ref. [24] which utilizes an IID security proof followed by the postselection technique lift [25, 26] to coherent attacks. The use of the postselection technique in Ref. [24] leads to pessimistic key rates, which is avoided by this work. Our proof of variable-length security uses the same essential tricks as prior work on variable-length security in Refs. [27] and [28, Chapter 3] for the phase error correction framework, and is nearly identical to [29, Supplementary Note A] for the EUR framework.
@@
This paper is organized as follows. In Section II, we describe the QKD protocol that we consider in this work. For simplicity, we first consider the BB84 protocol where Alice prepares perfect single-photon signal states. We do this since there are many existing techniques for dealing with imperfect state preparation by Alice, and the goal of this work is to focus on detector imperfections. We show that that the variable-length security of the QKD protocol follows if one is able to obtain suitable bounds on the phase error rate (Eq. 5). In Section III we show how such bounds can be obtained in the case where the basis-independent loss assumption is satisfied. In Section IV, we show how such bounds can be obtained in scenarios where the basis-independent loss assumption is not satisfied. We also discuss similarities and differences with other related work [11, 12, 15, 14, 13] towards the end of this section. In Section V we extend our analysis to prove the variable-length security of the decoy-state BB84 protocol with imperfect detectors. In Section VI we apply our results and compute key rates for the decoy-state BB84 protocol, and study the impact of basis-efficiency mismatch on key rates. We base our analysis of decoy-state BB84 on Lim et al [30], and also point out and fix a few technical issues in that work. In Section VII we explain how our results can be applied to detector side channels (Case. 3 above), and generalize to independent detectors. In Section VIII we summarize and present concluding remarks. Many details are relegated to the Appendices.
@@
    Resolve detector side-channel vulnerabilities by allowing Eve to control (bounded) detector imperfections.
@@
We have attempted to write this paper in a largely modular fashion beyond Section II. For variable-length security, one can directly read Appendices B and V.5. Similarly, for phase error estimation for the BB84 protocol, one can directly read Section III (for perfect detectors) and Section IV (for imperfect detectors). For phase error estimation in decoy-state BB84, one can read Sections V.2 and V.3. For a recipe for applying our results to compute key rates in the presence of detector imperfections, one can refer to Section VI.1. For the application of our results to detectors with efficiency mismatch, one can refer to Sections VI.2, VI.3 and VI.4.
@@ II Protocol Description
    where Alice’s POVMs are defined in Eq. 2, and Bob’s in Section VI.2.
@@
    This means that the probability of a round being inconclusive (i.e discarded) is independent of the basis. Notice that Eq. 4 depends on the choice of classical post-processing on Bob’s side. In particular, it can be trivially satisfied by mapping no-click and double-click events to 00 and $1$ randomly (so that $\Gamma^{(B)}_{(X,\bot)}=\Gamma^{(B)}_{(Z,\bot)}$ is zero). However, such a protocol cannot produce a key when loss is greater than $50\%$, and is therefore impractical. In general, if one assumes the canonical model of detectors (see Section VI.2), and maps double-clicks to 0/1 randomly, then Eq. 4 requires the loss and dark count rates in each detector-arm to be equal. This is why this condition is referred to as “basis-independent loss”, and its violation is referred to as “detection-efficiency mismatch” in the literature. Note that even for identical detectors, one is forced remap double-click events to satisfy Eq. 4.
@@
    Variable-Length Decision: When event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}$ occurs, Alice and Bob compute $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ (the number of bits to be used for one-way error-correction) and $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ (the length of the final key to be produced). Aborting is modelled as producing a key of length zero.
@@
    The variable-length protocol we consider allows users to determine the length of the error-correction information as a function of observations on the announced data. This data is anyway known to Eve, and therefore this procedure does not leak information via the above mechanism. Thus in this work, one must fix $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ to be a suitable function that determines the exact number of bits to be used for one-way error-correction, as a function of announcements. For more discussion, see footnote. 111 In general, the number of bits leaked during error-correction is equal to the length of the classical bit string needed to encode all possible transcripts of the error-correction protocol (which can be one-way or two-way). Thus, if one requires $\lambda_{\text{EC}}(...)$ to be an upper bound on the number of bits used, then we can proceed by noting that the number of bit strings of length up to some value $\lambda_{\text{EC}}$ is $12^{\lambda_{\text{EC}}+1}-$ 1, so a $(\lambda_{\text{EC}}+1)$-bit register suffices to encode all such bit strings. With this, it suffices to replace the $\lambda_{\text{EC}}(...)$ values in our subsequent key length formulas with $\lambda_{\text{EC}}(...)+1$. A similar analysis can be done for other error-correction protocols as well..
@@ II.1 Requirements on phase error estimation procedure and variable-length security
This states that the phase error rate is upper bounded (with high probability) by a suitable function $\mathcal{B}_{\delta_{1},\delta_{2}}$ of the observed error rate in the $X$ basis rounds, and the number of test and key generation rounds. We will obtain a suitable $\mathcal{B}_{\delta_{1},\delta_{2}}$ satisfying Eq. 5 in Sections III and IV, with and without the basis-independent loss assumption. The function $\mathcal{B}_{\delta_{1},\delta_{2}}$ depends on the metrics $\delta_{1},\delta_{2}$ that quantify the deviation from ideal behaviour for a given protocol description. We compute explicit bounds for $\delta_{1},\delta_{2}$ for detectors with efficiency mismatch in Section VI.3 using the recipe outlined in Section VI.1.
@@
Given an upper bound on the phase error rate (Eq. 5), we have the following theorem regarding the variable-length security of the QKD protocol described above. The proof is essentially identical to [29, Supplementary Note A], and uses the same techniques as those in Refs. [27, 28] and is included in Appendix B.
@@
where $h(x)$ is the binary entropy function for $x\leq 1/2$, and $h(x)=1$ otherwise. Then the variable-length QKD protocol that produces a key of length $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ using $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ bits for error-correction, upon the event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}\wedge\Omega_{\text{EV}}$ is $(2\varepsilon_{\text{AT}}+\varepsilon_{\text{PA}}+\varepsilon_{\text{EV}})$-secure 333 For pedagogical reasons, we ignore the issues arising from non-integer values of hash-lengths. Such issues can be easily fixed by suitable use of floor and ceiling functions..
@@ III.3 Sampling
$\Pr(\bm{e^{\text{key}}_{X}}\geq\bm{e^{\text{obs}}_{X}}+\gamma_{\text{serf}})_{|\Omega_{(n_{X},n_{K})}}$ $\leq e^{-\gamma_{\text{serf}}^{2}f_{\text{serf}}(n_{X},n_{K})},$ (12)
@@
$\gamma^{\varepsilon_{\text{AT}}}_{\text{serf}}(n_{X},n_{K})\coloneq\sqrt{\frac{\ln(1/\varepsilon_{\text{AT}}^{2})}{f_{\text{serf}}(n_{X},n_{K})}}\implies e^{-\left(\gamma^{\varepsilon_{\text{AT}}}_{\text{serf}}(n_{X},n_{K})\right)^{2}f_{\text{serf}}(n_{X},n_{K})}=\varepsilon_{\text{AT}}^{2}.$ (13)

### v3

@@ I Introduction
Security proofs of QKD based on the entropic uncertainty relations (EUR) [1, 2, 3], and the phase error correction approach [4, 5], yield some of the highest key rates against coherent attacks in the finite-size regime. While source imperfections [6, 7, 8, 9] have been extensively studied within the phase error correction framework, detector imperfections have not yet been addressed in any meaningful sense in either security proof framework. In particular, these proof techniques require the probability of detection in Bob’s detection setup to be independent of basis choice. This assumption is referred to as “basis-independent loss” in the literature, while the violation of this assumption is referred to as “basis-efficiency mismatch” or “detection-efficiency mismatch”. Satisfying this assumption requires the efficiency and dark count rates of Bob’s detectors to be exactly identical. Therefore, justifying this assumption in practice requires exact characterization of Bob’s identical detectors. Either of these tasks are impossible in practice. Therefore, these proof techniques are not applicable to practical prepare-and-measure (and entangled-based) QKD scenarios involving realistic detectors. Note that MDI-QKD [10] is able to address all detector imperfections and detector side-channels, since it assumes the detectors to be completely in Eve’s control. Morever, source imperfections can also be handled to a large degree. However, MDI-QKD is significantly more complex to implement as compared to prepare-and-measure QKD, and the latter remain dominant in real-world implementations. Thus, addressing detector imperfections within prepare-and-measure QKD is of paramout importance. To date, there is no security proof method that addresses this problem without requiring significant protocol modifications. While some theoretical analyses of basis-efficiency mismatch in the asymptotic setting exist for standard QKD [11, 12, 13, 14, 15], there is no work that handles basis-efficiency mismatch for coherent attacks in the finite-size regime. Meanwhile, there have been several experimental demonstrations [16, 17, 18, 19, 20], exploiting basis-efficiency mismatch for attacks on QKD systems.
In this work, we prove the security of the decoy-state BB84 protocol with an active detection setup without assuming basis-independent loss. We do so by showing that the phase error rate can be suitably bounded even without the assumption. We explicitly define metrics $\delta_{1}$, $\delta_{2}$ that quantify the deviation from the ideal case, and bound the phase error rate in terms of these deviations. Our framework is general, and can be applied to any (IID) detector model of one’s choice, as long as the relevant metrics $\delta_{1}$, $\delta_{2}$ can be suitable bounded. In general, this will require some model of the detectors and characterization of its properties. An important, non-trivial open question remains on how best to experimentally bound these quantities. We leave a full experimental analysis for future work, and restrict our attention in this work to the case where we utilise common models for detectors  [^1] We stress that characterisation will always proceed by assuming some physically motivated model of the devices in terms of a small number of paramters. The characterisation experiment will then estimate these model parameters. The question that must then be answered is about which model best describes the physical implementation, and how to characterise the parameters of the chosen model., and bound $\delta_{1},\delta_{2}$ in terms of commonly measured model parameters [22].
In this work we explicitly compute these metrics for the case of detectors with basis-efficiency mismatch and unequal dark count rates. To do so, we assume the the canonical model of detectors described in Section VI.3. The block-diagonal structure of the detector POVMs significantly aids the computation of these metrics. Moreover, we compute these metrics directly from the experimental characterization of the detection efficiencies and dark count rates of the detectors. Our results extend the security of QKD to the following practical scenarios:
@@
    Bob’s detectors are not identical, but the values of efficiency ($\eta_{b_{i}}$ for basis $b$ and outcome $i$) and dark count rates ($d_{b_{i}}$) are known. Note that while this is a useful toy model, such scenarios are impractical since they require $\eta_{b_{i}},d_{b_{i}}$ to be known exactly. We treat the dark count rate as a part of the POVM element, as described in Section VI.3.
@@
    Bob’s detectors are not identical, and the values of efficiency and dark count rates are only known to be in some range. Moreover, these values depend on the spatio-temporal modes (labelled by $\mathbf{d}$) of the incoming photons, and can therefore be chosen by Eve [17, 13, 19, 20]. This is expressed mathematically as $\eta_{b_{i}}(\mathbf{d})\in[\eta_{\text{det}}(\mathbf{d})(1-\Delta_{\eta}),\eta_{\text{det}}(\mathbf{d})(1+\Delta_{\eta})],d_{b_{i}}(\mathbf{d})\in[d_{\text{det}}(1-\Delta_{\text{dc}}),d_{\text{det}}(1+\Delta_{\text{dc}})]$. Note that in this model, the range of allowed values of the loss can depend on the spatio-temporal mode, whereas the dark count rates for all the modes lie in the same range.
@@
Our metrics $\delta_{1},\delta_{2}$ involve an optimization over all possible values of $\eta_{b_{i}},d_{b_{i}}$ in their respective ranges. Moreover, for our model of multi-mode detectors, we find that our methods yield the same values for Case 2 and Case 3. Thus, our methods address one practically important detector side-channel as a by-product. We discuss this in greater detail in Section VIII.
@@
We use entropic uncertainty relations [1] for our security proofs in this work. Since both the entropic uncertainty relations approach and the phase error correction [4, 5] approach involve bounding the phase error rate, we expect our techniques to also work when using the phase error correction approach for the security proof. In fact, one may also use the equivalence by Tsurumuru [23, 24] to relate the two security proof approaches. For example, [24, Section III] constructs an explicit phase error correction circuit whose failure probability can be bound by the smooth min entropy via [24, Theorem 1 and Corollary 2]. Therefore one may indirectly prove security in the phase error correction framework by applying the above equivalence on the results of this work.
@@
Moreover, we prove security in the variable-length framework [25, 26] which allows Alice and Bob to adaptively determine the number of bits to be used for error-correction, and the length of the output key, based on the observations during runtime. Such protocols are critical for practical implementations, where the honest behavior of the channel connecting Alice and Bob is difficult to determine in advance. A security proof for variable-length QKD has been recently obtained for generic protocols in Ref. [27] which utilizes an IID security proof followed by the postselection technique lift [28, 29] to coherent attacks. The use of the postselection technique in Ref. [27] leads to pessimistic key rates, which is avoided by this work. Our proof of variable-length security uses the same essential tricks as prior work on variable-length security in Refs. [30] and [31, Chapter 3] for the phase error correction framework, and is nearly identical to [32, Supplementary Note A] for the EUR framework.
@@
This paper is organized as follows. In Section II, we describe the QKD protocol that we consider in this work. For simplicity, we first consider the BB84 protocol where Alice prepares perfect single-photon signal states. We do this since there are many existing techniques for dealing with imperfect state preparation by Alice, and the goal of this work is to focus on detector imperfections. We show that that the variable-length security of the QKD protocol follows if one is able to obtain suitable bounds on the phase error rate (Eq. 5). In Section III we show how such bounds can be obtained in the case where the basis-independent loss assumption is satisfied. In Section IV, we show how such bounds can be obtained in scenarios where the basis-independent loss assumption is not satisfied. We also discuss similarities and differences with other related work [11, 33, 15, 14, 13, 34, 12] towards the end of this section. In Section V we extend our analysis to prove the variable-length security of the decoy-state BB84 protocol with imperfect detectors. In Section VI we apply our results and compute key rates for the decoy-state BB84 protocol, and study the impact of basis-efficiency mismatch on key rates. We base our analysis of decoy-state BB84 on Lim et al [35], and also point out and fix a few technical issues in that work. In Section VII we outline an approach towards addressing correlated detectors. In Section VIII we explain how our results can be applied to detector side channels (Case. 3 above). In Section IX we summarize and present concluding remarks. Many details are relegated to the Appendices.
@@
    Address some detector side-channel vulnerabilities by allowing Eve to control (bounded) detector imperfections.
@@
We have attempted to write this paper in a largely modular fashion beyond Section II. For variable-length security, one can directly read Appendices B and V.5. Similarly, for phase error estimation for the BB84 protocol, one can directly read Section III (for perfect detectors) and Section IV (for imperfect detectors). For phase error estimation in decoy-state BB84, one can read Sections V.2 and V.3. For a recipe for applying our results to compute key rates in the presence of detector imperfections, one can refer to Section VI.1. For the application of our results to detectors with efficiency mismatch, one can refer to Sections VI.3, VI.4 and VI.5. Sections VII and VIII can be read independently, but require Section IV to be read first.
@@ II Protocol Description
    where Alice’s POVMs are defined in Eq. 2, and Bob’s in Section VI.3.
@@
    This means that the probability of a round being inconclusive (i.e discarded) is independent of the basis for all input states. Notice that Eq. 4 depends on the choice of classical post-processing on Bob’s side. In particular, it can be trivially satisfied by mapping no-click and double-click events to 00 and $1$ randomly (so that $\Gamma^{(B)}_{(X,\bot)}=\Gamma^{(B)}_{(Z,\bot)}$ is zero). However, such a protocol cannot produce a key when loss is greater than $50\%$, and is therefore impractical. In general, if one assumes the canonical model of detectors (see Section VI.3), and maps double-clicks to 0/1 randomly, then Eq. 4 requires the loss and dark count rates in each detector-arm to be equal. This is why this condition is referred to as “basis-independent loss”, and its violation is referred to as “detection-efficiency mismatch” in the literature. Note that even for identical detectors, one is forced remap double-click events to satisfy Eq. 4.
@@
    Variable-Length Decision: When event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}$ occurs, Alice and Bob compute $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ (the number of bits to be used for one-way error-correction) and $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ (the length of the final key to be produced). Aborting is modeled as producing a key of length zero.
@@
    The variable-length protocol we consider allows users to determine the length of the error-correction information as a function of observations on the announced data. This data is anyway known to Eve, and therefore this procedure does not leak information via the above mechanism. Thus in this work, one must fix $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ to be a suitable function that determines the exact number of bits to be used for one-way error-correction, as a function of announcements. For more discussion, see footnote. 222 In general, the number of bits leaked during error-correction is equal to the length of the classical bit string needed to encode all possible transcripts of the error-correction protocol (which can be one-way or two-way). Thus, if one requires $\lambda_{\text{EC}}(...)$ to be an upper bound on the number of bits used, then we can proceed by noting that the number of bit strings of length up to some value $\lambda_{\text{EC}}$ is $12^{\lambda_{\text{EC}}+1}-$ 1, so a $(\lambda_{\text{EC}}+1)$-bit register suffices to encode all such bit strings. With this, it suffices to replace the $\lambda_{\text{EC}}(...)$ values in our subsequent key length formulas with $\lambda_{\text{EC}}(...)+1$. A similar analysis can be done for other error-correction protocols as well..
@@ II.1 Requirements on phase error estimation procedure and variable-length security
This states that the phase error rate is upper bounded (with high probability) by a suitable function $\mathcal{B}_{\delta_{1},\delta_{2}}$ of the observed error rate in the $X$ basis rounds, and the number of test and key generation rounds. We will obtain a suitable $\mathcal{B}_{\delta_{1},\delta_{2}}$ satisfying Eq. 5 in Sections III and IV, with and without the basis-independent loss assumption. The function $\mathcal{B}_{\delta_{1},\delta_{2}}$ depends on the metrics $\delta_{1},\delta_{2}$ that quantify the deviation from ideal behavior for a given protocol description. We compute explicit bounds for $\delta_{1},\delta_{2}$ for detectors with efficiency mismatch in Section VI.4 using the recipe outlined in Section VI.1.
@@
Given an upper bound on the phase error rate (Eq. 5), we have the following theorem regarding the variable-length security of the QKD protocol described above. The proof is essentially identical to [32, Supplementary Note A], and uses the same techniques as those in Refs. [30, 31] and is included in Appendix B.
@@
where $h(x)$ is the binary entropy function for $x\leq 1/2$, and $h(x)=1$ otherwise. Then the variable-length QKD protocol that produces a key of length $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ using $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ bits for error-correction, upon the event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}\wedge\Omega_{\text{EV}}$ is $(2\varepsilon_{\text{AT}}+\varepsilon_{\text{PA}}+\varepsilon_{\text{EV}})$-secure 444 For pedagogical reasons, we ignore the issues arising from non-integer values of hash-lengths. Such issues can be easily fixed by suitable use of floor and ceiling functions..
@@ III.3 Sampling
$\Pr(\bm{e^{\text{key}}_{X}}\geq\bm{e^{\text{obs}}_{X}}+\gamma_{\text{serf}})_{|\Omega_{(n_{X},n_{K})}}$ $\leq e^{-2\gamma_{\text{serf}}^{2}f_{\text{serf}}(n_{X},n_{K})},$ (12)
@@
$\gamma^{\varepsilon_{\text{AT}}}_{\text{serf}}(n_{X},n_{K})\coloneq\sqrt{\frac{\ln(1/\varepsilon_{\text{AT}}^{2})}{2f_{\text{serf}}(n_{X},n_{K})}}\implies e^{-2\left(\gamma^{\varepsilon_{\text{AT}}}_{\text{serf}}(n_{X},n_{K})\right)^{2}f_{\text{serf}}(n_{X},n_{K})}=\varepsilon_{\text{AT}}^{2}.$ (13)